OpenStudy (anonymous):

4cos x-4 sin ^2 x+5=0

OpenStudy (anonymous):

\[0 \le x \le \pi\]

OpenStudy (raden):

convert sin^2 x to 1-cos^2 x, to get the cos all terms

OpenStudy (anonymous):

so then it'd be 4cos x-4-4cos^2x?

OpenStudy (raden):

4cosx - 4 sin^2 x + 5 = 0 4cosx - 4(1-cos^2 x) + 5 = 0 4cosx - 4 + 4cos^2 x + 5 = 0 4cos^2 x + 4cosx + 1 = 0

OpenStudy (raden):

factor out this, (2cosx + 1)(2cosx + 1) = 0

OpenStudy (raden):

find the zeroes by take the factors be zero 2cosx + 1 = 0

OpenStudy (anonymous):

so i split the 2cosx and the 1 and solve for both?

OpenStudy (raden):

2cosx + 1 = 0 2cosx = -1 cosx = -1/2 yes, solve for x's

OpenStudy (anonymous):

The answer to the question is 2pi/3

OpenStudy (raden):

yeah, in interval 0≤x≤π the solution only satisfied for x=2pi/3

OpenStudy (anonymous):

so how did we go from -1/2 to 2pi/3? I'm missing a mental step somewhere

OpenStudy (raden):

u have to memorize the values fo sine, cosine, etc... for the special angles ::)

OpenStudy (anonymous):

Well yuck. Thanks!